slope=3 over 4 and a point (-2,6)

Answer by mananth(16946)

slope=3 over 4 and a point (-2,6)
m= 3/4 (-2,6)
..
find b in y = mx +b
plug the values
6=-2(3/4)+b
6=-3/2+b
6+3/2= b
b=15/2
...
y= 3x/4 +15/2
..
m.ananth@hotmail.ca